A 3000-Kg is hoisted 2 m above the ground. Calculate its potential energy
1
Expert's answer
2019-03-19T12:20:47-0400

We can find the potential energy from the formula:


"PE = mgh = 3000 kg \cdot 9.8 \dfrac{m}{s^2} \cdot 2 m = 5.88 \cdot 10^4 J."

Answer:

"PE = 5.88 \cdot 10^4 J."